OpenStudy (anonymous):

how to find the x-intercepts of (5x-3)(x+7)/(-3x-1)(7x-3)

OpenStudy (anonymous):

Hm let's see Assuming I'm reading this right: \[ f(x) = \frac{(5x-3)(x+7)}{(-3x-1)(7x-3)} \] Well first, we know that the divisor can't be 0.. so \[ (-3x-1)(7x-3) \ne 0 \quad \text{ means:}\\ -3x-1\ne 0 \quad \text{and} \quad 7x-3 \ne 0 \\ x \ne -\frac{1}{3} \quad \text{and} \quad x \ne \frac{3}{7} \] Now we say: \[ f(x) = 0 \quad \implies \quad \frac{(5x-3)(x+7)}{(-3x-1)(7x-3)} = 0 \] For f(x) to be 0 the numerator has to be 0.. (or simply multiply both side by the divisor..) \[ (5x-3)(x+7) = 0 \quad \text{ means:}\\ 5x-3 = 0 \quad \text{or} \quad x+7 = 0 \\ x = \frac{3}{5} \quad \text{or} \quad x = -7 \\ \] Since both results are legal (won't make a division by 0), then we can say, the function has value 0 (intersects X axis) when x =3/5 or x = -7

OpenStudy (anonymous):

thank you!